OpenStudy (anonymous):

i need help with this problem 2 (3x + 4) = x + 6

OpenStudy (anonymous):

1)The objective is to get "X" by itself. 2) On the left side of the equation, multiply the 2 by both numbers inside the parenthesis. (6x+ 12) 3) Subtract the 6x from the both sides, so the right side is -5x. x- 6x= -5x 4) Subtract the -6 on each side of the equation so that there is only -5x on the right side. 5) Use the numbers you have solved for on the left side and divide by -5 to get the variable "x" by itself.

OpenStudy (anonymous):

dude i think you did it wrong

OpenStudy (anonymous):

2 x 4 is 8

OpenStudy (anonymous):

Yes. That is where you come in. :) I only explained how to get the x by itself. Finding the number on the left (8) was up to you. So now you have 8 just got to step 4/

OpenStudy (anonymous):

go to* 6x+8= x+6 > 6x+2=x> 2= -5x -2/5=x Is this the answer you got?